Check whether the limit of the function
f(x,y)=3x^3y/x^6+2y^2 exists as (x, y) →(0,0)
1
Expert's answer
2020-03-30T09:55:30-0400

"\lim_{x\to\ 0\\y\to\ 0}f(x,y)= \frac{3\frac{y}{x^3}}{1+2(\frac{y}{x^3})^2}= \frac{3k}{1+2k^2}" where "k=\frac{y}{x^3}" .The limit "f(x,y)" depends on "k" hence "f(x,y)" has no limit.